ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 49, 93 Gabasov Rafail, Kirillova Faina, Balashevich Nataliya, Institute for Mathematics of NAS, Belarus Synthesis of optimal dynamical systems The problem of constructing optimal state and output feedbacks is under consideration (D. W. Bushaw, 1953). A new approach to the problem based on introducing defining elements and defining (algebraic) equations is justified. A method of real-time solving the defining equations is suggested. The approach is applied to real-time construction of output feedback for control systems under uncertainty I(u) = h′_{0}x(t^{∗}) → max, ẋ = Ax + Bu + w(t), with a priori information x(0) ∈ X_{0}, w(t) ∈ W, x(t^{∗}) ∈ X^{∗} and a sensor y(t) = Cx(t) + ξ(t), ξ_{∗} ≤ ξ(t) ≤ ξ^{∗}, t ∈ [0, t^{∗}]. Results are used to investigate the stabilization problem. Reference: R. Gabasov, F. M. Kirillova, and S. V. Prischepova, Optimal Feedback Control, Lecture Notes in Control and Information Sciences, 207, Springer-Verlag, 1995.